the value v(t) insents of a share is give by v(t) = 2t^2-16t+40 where t represents the number of weeks since the shares purchase
a) after how many weeks is the share worth 58 cents?
b)can the share reach a value of 6 cents? justify your answerr

Answer by nerdybill(7384)

the value v(t) insents of a share is give by v(t) = 2t^2-16t+40 where t represents the number of weeks since the shares purchase
a) after how many weeks is the share worth 58 cents?
set v(t) to 58 and solve for t:
v(t) = 2t^2-16t+40
58 = 2t^2-16t+40
dividing both sides by 2:
29 = t^2-8t+20
0 = t^2-8t-9
factoring:
0 = (t+1)(t-9)
t = {-1, 9}
toss out the negative solution leaving:
t = 9 weeks
.
b)can the share reach a value of 6 cents? justify your answerr
answer: NO
set v(t) to 6 and solve for t:
v(t) = 2t^2-16t+40
6 = 2t^2-16t+40
3 = t^2-8t+20
0 = t^2-8t+17
calculate discriminate:
b^2 -4ac
64 - 4(1)(17)
64 - 68
-4
since the discriminate is negative there is NO real solutions